# Quality-Aware Translation Models: Efficient Generation and Quality Estimation in a Single Model

### v3

@@ 7 Conclusion
This paper introduces a novel approach to enhance NMT by making the models quality-aware. Our approach addresses the issue of misalignment between outputs generated via MAP decoding and human judgment. We achieve this by training NMT models to assess the quality of their own translations, effectively circumventing the limitations of conventional decoding methods. As a result this new approach yields improvements similar or superior to QE-reranking approaches, but with the efficiency of MAP-decoding, i.e. with single-pass decoding (for QA-prompting, the best performing method). QE-reranking in contrast needs a sampling step followed by a reranking step using an external, computationally expensive model.
@@
A potential risk of our method might be that it is more resource-intensive than simple models, and thus might increase the quality difference with respect to low-resource languages, since they are unlikely to be allocated as many resources as the high-resource languages.
@@ Appendix A Additional Results for Language Pair: English to Japanese
Baseline Full 76.3 85.7
Quality-Aware Prompting (Ours) Full 80.3 87.8
Quality-Aware Prediction (Ours) Full 80.9 87.7
Baseline Filter 77.8 86.2
Quality-Aware Prompting (Ours) Filter 80.0 87.7
Quality-Aware Prediction (Ours) Filter 79.6 87.5
@@ Appendix B BLEU scores
Baseline Full 80.2 85.8 35.4 1.81
QA Prompting (Ours) Full 82.3 87.1  ${}^{\star}$ 36.6 1.43  ${}^{\star}$
QA Prediction (Ours) Full 82.0 86.5  ${}^{\star}$ 27.1 2.07
External QE-Reranking Full 83.3 86.9  ${}^{\star}$ 28.1 1.50  ${}^{\star}$
Baseline Filter 81.8 87.0 36.2
QA Prompting (Ours) Filter 82.6 87.3  ${}^{\star}$ 36.7
QA Prediction (Ours) Filter 82.5 86.7 28.1
External QE-Reranking Filter 83.7 86.9 27.8
@@
External QE-Reranking 83.3 86.9  ${}^{\star}$ 28.1 1.50  ${}^{\star}$
$M=50$ MBR Baseline 82.5 86.8  ${}^{\star}$ 30.8 1.41  ${}^{\star}$
MBR QA Prompting (Ours) 83.5 87.4   ${}^{\star{\dagger}}$ 31.9 1.36   ${}^{\star{\dagger}}$
MBR QA Prediction (Ours) 83.4 87.5   ${}^{\star{\dagger}}$ 31.1 1.45   ${}^{\star{\dagger}}$
@@
MBR QA Prompting (Ours) 82.6 86.9  ${}^{\star}$ 31.3
MBR QA Prediction (Ours) 82.8 87.0  ${}^{\star}$ 29.1
@@
MBR QA Prompting (Ours) 83.8 87.7   ${}^{\star{\dagger}}$ 33.0
MBR QA Prediction (Ours) 83.9 87.7   ${}^{\star{\dagger}}$ 32.0
@@
MBR QA Prompting (Ours) 83.1 87.3  ${}^{\star}$ 32.4
MBR QA Prediction (Ours) 83.1 87.2  ${}^{\star}$ 30.2
@@ Appendix C Quality-Aware LLMs
In this section we explore the potential of enhancing LLMs with quality awareness through our proposed method. Specifically, we employ our most efficient approach, Quality-Aware Prompting, to finetune PaLM-2 Bison Anil et al. (2023). To this end, we finetune the pretrained model for 10k steps using our QA Prompting approach. As baseline we also finetune the LLM with identical configurations on the same data, but without incorporating any quality signal. In Table 7, we observe that QA Prompting substantially outperforms standard finetuning by 1.3 Bleurt points. Our results suggest that substantial performance improvements can be achieved with minimal data and steps. This opens the door to the possibility of utilizing more costly QE methods or even human evaluations in the future to curate finetuning datasets and align models directly with human preferences.
@@ Appendix E Influence of Bin Identifier Choice
Appendix F Example Translations for QA-prompting
@@
Source
Needless to say, it was chaos.
QA Label [0]
@@
Reference
Es war natürlich ein Chaos.
Source
The total number of vaccines injected in the country reached 109,990,742 doses.
QA Label [0]
@@
Reference
Die Gesamtzahl der im Land verabreichten Impfungen beläuft sich auf 109.990.742 Dosen.
Source
Long story short, she comes back to the checkout, but goes to another line.
QA Label [0]
@@
Reference
Um es kurz zu machen, sie kam zurück zur Kasse, stellte sich aber bei einer anderen Schlange an.
@@ Appendix H Details of Quality-Aware Approaches
In the context of Quality-Aware Translation, two approaches are explored. In Quality-Aware Prompting (QA Prompting), a quality score is appended to the source segment during training, allowing the model to associate quality score strings with examples of translation exhibiting the same level of quality (Figure 6). Since this quality token appears in the input, it allows for direct prompting of high-quality translations during decoding. Conversely, Quality-Aware Prediction (QA Prediction) involves training a model that predicts both a hypothesis and a quality score concurrently. This approach transforms the translation model into a Quality Estimation (QE) model by appending the quality score string to the target sentence during training, enabling the model to predict quality during inference (Figure 6).
[…]

### v4

@@ 7 Conclusion
This paper introduces a novel approach to enhance NMT by making the models quality-aware. Our approach addresses the issue of misalignment between outputs generated via MAP decoding and human judgment. We achieve this by training NMT models to assess the quality of their own translations, effectively circumventing the limitations of conventional decoding methods. As a result this new approach yields improvements similar or superior to QE-reranking approaches, but with the efficiency of MAP-decoding, i.e. with single-pass decoding (for QA Prompting, the best performing method). QE-reranking in contrast needs a sampling step followed by a reranking step using an external, computationally expensive model.
@@
A potential risk of our method might be that training it is more resource-intensive than simple models, and thus might increase the quality difference with respect to low-resource languages, since they are unlikely to be allocated as many resources as the high-resource languages.
@@ Appendix A Additional Results for Language Pair: English to Japanese
Baseline Full Corpus 76.3 85.7
Quality-Aware Prompting (Ours) Full Corpus 80.3 87.8
Quality-Aware Prediction (Ours) Full Corpus 80.9 87.7
Baseline Filtered 77.8 86.2
Quality-Aware Prompting (Ours) Filtered 80.0 87.7
Quality-Aware Prediction (Ours) Filtered 79.6 87.5
@@ Appendix B BLEU scores
Baseline Full Corpus 80.2 85.8 35.4 1.81
QA Prompting (Ours) Full Corpus 82.3 87.1   36.6 1.43
QA Prediction (Ours) Full Corpus 82.0 86.5   27.1 2.07
External QE-Reranking Full Corpus 83.3 86.9   28.1 1.50
Baseline Filtered 81.8 87.0 36.2
QA Prompting (Ours) Filtered 82.6 87.3   36.7
QA Prediction (Ours) Filtered 82.5 86.7 28.1
External QE-Reranking Filtered 83.7 86.9 27.8
@@
External QE-Reranking 83.3 86.9   28.1 1.50
$M=50$ MBR Baseline 82.5 86.8   30.8 1.41
MBR QA Prompting (Ours) 83.5 87.4   ⋆† 31.9 1.36   ⋆†
MBR QA Prediction (Ours) 83.4 87.5   ⋆† 31.1 1.45   ⋆†
@@
MBR QA Prompting (Ours) 82.6 86.9   31.3
MBR QA Prediction (Ours) 82.8 87.0   29.1
@@
MBR QA Prompting (Ours) 83.8 87.7   ⋆† 33.0
MBR QA Prediction (Ours) 83.9 87.7   ⋆† 32.0
@@
MBR QA Prompting (Ours) 83.1 87.3   32.4
MBR QA Prediction (Ours) 83.1 87.2   30.2
@@ Appendix C Quality-Aware LLMs
In this section we explore the potential of enhancing LLMs with quality awareness through our proposed method. Specifically, we employ our most efficient approach, Quality-Aware Prompting, to finetune PaLM-2 Bison Anil et al. (2023). To this end, we finetune the pretrained model for 10k steps with the same data as in the main text (en-de) using our QA Prompting approach. As baseline we also finetune the LLM with identical configurations on the same data, but without incorporating any quality signal. In Table 7, we observe that QA Prompting substantially outperforms standard finetuning by 1.3 Bleurt points. Our results suggest that substantial performance improvements can be achieved with minimal data and steps. This opens the door to the possibility of utilizing more costly QE methods or even human evaluations in the future to curate finetuning datasets and align models directly with human preferences.
@@ Appendix E Influence of Bin Identifier Choice
Appendix F Example Translations for QA Prompting
@@
Source Needless to say, it was chaos.
@@
Reference Es war natürlich ein Chaos.
Source The total number of vaccines injected in the country reached 109,990,742 doses.
@@
Reference Die Gesamtzahl der im Land verabreichten Impfungen beläuft sich auf 109.990.742 Dosen.
Source Long story short, she comes back to the checkout, but goes to another line.
@@
Reference Um es kurz zu machen, sie kam zurück zur Kasse, stellte sich aber bei einer anderen Schlange an.
@@ Appendix H Details of Quality-Aware Approaches
In the context of quality-aware translation, two approaches are explored. In Quality-Aware Prompting (QA Prompting), a quality score is appended to the source segment during training, allowing the model to associate quality score strings with examples of translation exhibiting the same level of quality (Figure 6(a)). Since this quality token appears in the input, it allows for direct prompting of high-quality translations during decoding. Conversely, Quality-Aware Prediction (QA Prediction) involves training a model that predicts both a hypothesis and a quality score concurrently. This approach transforms the translation model into a Quality Estimation (QE) model by appending the quality score string to the target sentence during training, enabling the model to predict quality during inference (Figure 6(b)).
(a) Quality-Aware Prompting
(b) Quality-Aware Prediction